OpenStudy (anonymous):

Lyle has a ten-cup box of cereal. If one serving is three-fourths cup, how many servings are in the box? thirteen and one-third 7 and one-half twelve and two-thirds ten and three-fourths

OpenStudy (anonymous):

7 and a half?

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (anonymous):

@student_basil

OpenStudy (anonymous):

its 7 1/2 correct?

OpenStudy (anonymous):

yes its 7 1/2

OpenStudy (anonymous):

you know that there is an amount of 10 cups ... each serving is 3/4 of a cup... So the number of servings would be 10 ÷ (3/4) 10 ÷ (3/4)= 10 × (4/3)= 40/3= (39+1) /3= 13 1 /3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

Yes :))

OpenStudy (anonymous):

@YanaSidlinskiy is he correct

OpenStudy (yanasidlinskiy):

Why are you doubting??

OpenStudy (anonymous):

10 times 3/4 = 7.5

OpenStudy (yanasidlinskiy):

Yes! @student_basil is correct!

OpenStudy (anonymous):

ok cool

OpenStudy (yanasidlinskiy):

No problem!